OpenStudy (anonymous):

Write the 6 trig values in exact form for the acute angle formed by the line passing through the point: P(3, 4)

OpenStudy (selrachcw95):

I am looking. :D

OpenStudy (selrachcw95):

Not sure if this is what they want but here is what I found sin θ = 5/4, cos θ = 3/5, tan θ = 4/3

OpenStudy (cwrw238):



OpenStudy (anonymous):

i figured it out :) thank you guys for helping

OpenStudy (cwrw238):

sin = 4/5 , cos = 3/5 and tan = 4/3

OpenStudy (selrachcw95):

Was I right?

OpenStudy (selrachcw95):

Or even close to right?

OpenStudy (cwrw238):

you had sin and cos wrong the point (3,4) means that adjacent(x) length is 3 and opposite side ( y) = 4